OpenStudy (decentnabeel):

if there are 3 dogs for every 2 cats is an animal shelter and the shelter has a total of 900 animals how many are dogs?

OpenStudy (decentnabeel):

@Directrix

OpenStudy (faiqraees):

\[\large\rm Number ~of~ dogs = \frac{Number ~of~ dogs}{Number ~of~ dogs~and~cats}*900 \]

Directrix (directrix):

The ratio of dogs to cats is 3/2. That comes from the given info that for every 3 dogs, there are 2 cats. D/C = 3/2 where D is the number of dogs and C is the number of cats 3*C = 2*D C = 2/3 * D That gives an equation for you to solve. (2*D) / 3 + D = 900 @DecentNabeel

Directrix (directrix):

@DecentNabeel Solve for D, the dogs.